OpenStudy (ladylauren):

Simplify 4/5n-1/5=2/5n

OpenStudy (jiteshmeghwal9):

collect all the like terms on one side \[\frac{4}{5}n-\frac{2}{5}n=\frac{1}{5}\]

OpenStudy (jiteshmeghwal9):

can u solve it further

OpenStudy (ladylauren):

2/5n=1/5